(f)In order to find the height of a bridge, a rock was dropped off the bridge and timed.
If it took three seconds to hit the water, what is the height of the bridge? 2marks
4. (a) Determine the distance covered by a body in 4.0s if its initial velocity, u is 20 m/s
and acceleration is 4.0 m/s2
. (4
marks)
(b) state three effects of forces on objects (6 marks)
(c) State two demerits of Friction and suggest two ways of reducing it. (4marks)
(d) Explain three Clinical Applications of Gravity (
Answer #1

F). Using the equation of motion in this situation

v= u + gt

v= 0 + 9.8 x 3 = 29.4 m/ s

For height of the bridge , use

v^2 - u^2 = 2gh

( 29.4)^2 = 2 x 9.8 x h

h = 44.8 m

4) (a). Using. s= ut + 1/2 at^2

s = 20 x 4 +1/2 x 4 x4^2

s= 112m

b) Force can make the body in motion or stop.

It can change the shape and size .

It can change the direction of the body.

c ) when a body is in motion including the friction then it causes the body moving parts shear n tear.

It produces heat.

For both preventing the heat production n wear n tear of moving part,. Lubrication is generally used to reduce the friction.

Use of rolling elements bearing are used.

d) It causes to come back of a ball with some velocity when it is thrown upwards.

It makes the moon to revolve around the earth.

It makes every one to walk on the earth instead of floating on the space.